OpenStudy (anonymous):

rope pulling a sled is sloped at 25° upward at the point of contact. The tension in the rope is 40.0 N. What is the horizontal component of the force?

OpenStudy (mos1635):

 Fx=F*cosΘ Fx= 40*cos25 Fx=40*0,9 Fx=36N

OpenStudy (anonymous):

 here, F(horizontal)=tension in rope*cosΘ F(horizontal)=40*cos25 so, F(horizontal)=36N ans